OpenStudy (anonymous):

f(x)= 4x+3/5x+2 find f ' (x)

OpenStudy (anonymous):

\[(5x+2(4)-(4x+3)(5))/ (5x+2)^2\]

OpenStudy (anonymous):

\[5x+8-(20x+15)\]/(5x+2)^2

OpenStudy (anonymous):

\[-15x-5/(5x+2)^2\]

OpenStudy (anonymous):

sorry -7 not -5

OpenStudy (anonymous):

use \[(\frac{f}{g})'=\frac{gf'-fg'}{g^2}\] with \[f(x)=4x+3,f'(x)=4,g(x)=5x+2,g'(x)=5\] and you will get it

OpenStudy (anonymous):

can that reduce?

OpenStudy (anonymous):

don't look to reduce. it is rare

OpenStudy (anonymous):

just multiply out and combine like terms in the numerator

OpenStudy (anonymous):

f'(x) =-7/(5x+2)^2

OpenStudy (anonymous):

\[4x+\frac{3 x}{5}+2=2+\frac{23 x}{5} \]Derivative of 2+23*x/5 = 23/5